write the slope intercept equation for the line that passes through (2,-2)(-1,10)
Answer by richwmiller(17219)

-2=2m+b
10=-m+b
b=10+m
-2=2m+10+m
-12=3m
-4=m
10=4+b
6=b
y=-4x+6